OpenStudy (anonymous):

I WILL FAN AND MEDAL, 8 times the square of a nonzero number is equal to 96 times the number.What is the number? A)12 B)24 C)96 D)4

OpenStudy (anonymous):

8X^2 = 96X

OpenStudy (anonymous):

A)12

OpenStudy (anonymous):

can u explain @Brian42 and @Ruby-Gloom

OpenStudy (anonymous):

Yeah 8 x X = 92 Replace X with your answers until you get 92.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

8x^2 - 96x = x(8x-96) factor 8x = 96 Solve for x x=12